• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Or Chen

    Log entry time 00:08:23 on June 9,2010

    Entry number 325340

    keyword=Summary

    Swing Shift Starting 16:00 - Continue prodution.
    SL Or Chen
    TO conrad aniol
    3rd kiadtisak Saenboonruang




    16:00 - Shift Started.
    Curent Status: Taking data.

    16:10 - Compton DAQ lost conection to ROC-6.
    Run Number: 22984.
    Hardware rebooted. Problem sloved.

    19:15 - R-HRS alerts D-N D-G P0 value is: 1.07016 GeV/c
    Stoped run.

    19:17 - mcc reported vacum problem.
    No beam for now.

    19:42 - Start new run: 4413
    NOTE: There is still a problem with the right HRS.

    19:54 - Stop Run.
    Stop beam.
    Request MMC for escorted acces.

    20:30 - Jack Segal reported that he can't fix the R-HRS dipole magnet.
    Will continue runing using only the L-HRS.

    20:42 - Beam restored, Run number: 4414.

    23:30 - Alarm started for no clear reason.
    See dedicated log entry,
    Next SL took over.

    24:00 - Shift over.
    Current Status: Taking data with no R-HRS